# Monte Carlos for tau lepton – Standard Model

### v3

@@ Abstract
One of purposes for High Energy accelerator experiments is confrontation of theory and measurements in ever new realms. Any new agreement extends theory applicability domain, any discrepancy hints to unexplained. That calls for better calculations or new, deeper theory. Often one has to search for small contributions over large Standard Model background. Multi-dimensional signatures and complex background subtraction cuts imply that Monte Carlo techniques are indispensable. My talk included description of: KKMC Monte Carlo for $e^{+}e^{-}\to\tau^{+}\tau^{-}(n\gamma)$ (with $\tau$ decays), generation of additional lepton pairs of SM and New Physics, and contributions from anomalous magnetic and electric dipole moments.
@@ 1 Introduction
My talk is devoted to KKMC Precision Monte Carlo for $e^{+}e^{-}\to\tau^{+}\tau^{-}(n\gamma)$ , with $\tau$ decays included, and its associated projects. The main design purposes, as in the past, include: (i) precision simulation of lepton pair production, (ii) flexible set up for initialization of user defined hadronic currents used in $\tau$ decays. There are new aspects which gain importance. The C++ version of KKMC is available, which has better than the FORTRAN predecessor arrangements for beam-spread for example. But the most important aspect was to preserve continuity of numerical tests, essential ingredient of work see Ref. [1]. The long road toward enhanced precision of QED part started already  [2], but this road toward third order matrix elements is long and perilous. There is a need for new contributors to enter, to preserve skills and continuity of the projects, and at the same time to address new challenges.
Let us now turn attention to specific projects, to their limitations too. The C++ version of KKMC is public, but is not yet adopted to lower energies (the function for photon vacuum polarization $\Pi_{\gamma\gamma}(s)$ to be obtained from dispersion relations is clearly defined, but needs to be then adjusted). There are clear manpower issues: help in installations, set ups and of some phenomenology applications is not as robust as it was in the past. For the F77 version as installed in Belle2 I take care and help, but main burden is now on somebody’s else shoulders.
Let us turn now attention to generation of additional lepton pairs originating from SM or New Physics interactions. Pair emission is a necessary step toward third order matrix element installation. From the algorithmic point of view, treated as radiative corrections it is incompatible with exponentiation, because corresponding crude distribution does not feature conformal symmetry. This symmetry is essential in building relation between crude level Monte Carlo with matrix elements of eikonal form. It is also important for matching line-shape distribution (intermediate resonances) and initial state bremsstrahlung. Alternative solutions are necessary. Real pair emission can be achieved by running of KKMC with appropriate flags for pair emission part of loop vertex correction switched on. Then simultaneous run, of four fermion final state Monte Carlo program, like KORALW [3], is necessary. Alternatively, for final state pair emissions, Monte Carlo of the after-burn type, PHOTOS is available [4]. Program enables full phase-space coverage. Although tests with Matrix element simulations indicate that precision is sufficient for today applications, some studies of related systematic ambiguities are included in Ref. [5]. These publications do not match the FCC precision, and further work is needed.
@@
Let us now turn attention to more technical aspects. The process of migration from FORTRAN to C++ is advancing well. The C++ Versions of PHOTOS Monte Carlo for radiative corrections in decays [4] and KKMC [1] are already published. The TAUOLA Monte Carlo for $\tau$ decays is prepared for such transformation too, but actual implementation is not completed. It’s internal structure was already prepared some time ago [10]. Further improvements are now in Belle 2 hands. Fits and evaluation of ambiguities of explored data are necessary. Program was prepared to run with user provided hadronic currents, which could be coded in C++ and activated with redefinition of pointers. That solution was temporarily but partly abandoned. Step back was due to requirements of Belle software organization.
My experience with PHOTOS where migration to C++ has been completed completed is that the best is to translate program with young researchers participation. They can observe that migration does not go too fast, because existing arrangements for future and for tests could be then easily lost. Participation of researchers with distinct profile is helpful, and prepares us for future takeover of the projects, independently whoever it may be.
The software for Belle 2 is working well and I am involved personally. Important contributor is now Swagato Banerjee. Installation and initialization for other platforms pose now-days problems especially for C++ version of KKMC. For example proper loading of root library was bringing difficulties and it was reported to us but our response was slow. That is one of the most visible aspect of the man power issue for our community, but passing expertise to the next generation is potentially larger issue, even if related challenges are not imminent to see.
@@ 2 Basics of precision Monte Carlo
Let me stress at first KKMC is based on “matrix element $\times$ full and exact phase space”. Neither KKMC nor PHOTOS employ shower-like algorithm. They are both non-Markovian. Generation starts with Poisson distribution in number of photon candidates. In case of KKMC tower of simplified/improvable structures starts from one-dimensional spectrum, then internal generator exploiting eikonal matrix elements valid all over full multi-photon phase space is devised. It was highly non-trivial step to achieve. Next step is exploiting bremsstrahlung part of QED matrix elements and Yennie–Frautschi–Suura exponentiation [11]. That is not all what is needed: the electroweak and strong interaction parts need to be taken in as well. Finally, light lepton pair emissions must be taken into account. I have omitted complexity of presamplers necessary due to collinear configurations, which are naturally regulated by electron (or other lepton) masses. Finally library of Matrix Elements devised as input for event weight, sometimes called “model weight”, represents independent module see Fig. 1. Term “model weight” may be misunderstood. There is no room for arbitrary modeling. Amplitudes in use are the result of formal fixed order matrix elements. For details see [12].
@@
Note that for phase space organization, in KKMC design, conformal symmetry of multi-photon phase space and eikonal parts of QED amplitudes is fundamental. That is why, algorithm is suitable for many applications, path for implementation fixed order amplitudes is defined. At present, second order matrix elements are installed. In case of PHOTOS different solution is used for exact multi-photon phase space generation. It is restricted to final state radiations only, and solution is not explored beyond use of the first order matrix elements. On the other hand, emission of massive particles was easier to implement. For details seem [5, 4] and references therein.
@@ 3 Additional pairs
For pair emissions in final states modified algorithm of PHOTOS Monte Carlo can be used. In this case phase space is treated exactly and no approximations are used, independently of whether the pair emission is based on QED calculated matrix element or of New Physics model. As effects of pair emissions are not large, matrix elements are simplified (in improvable way though)! Tests/comparisons with matrix element based calculations are essential part of the work. As an example Fig.  2 is provided.
@@ 4 Anomalous dipole moments
For the simulation of New Physics effects, simplified kinematic for implementation is usually sufficient. There can be two weights calculated, for the cross section:
@@
The $R_{ij}$ depends on kinematic of $\tau$-pair production, $h^{i}_{\pm}$ on $\tau^{\pm}$ decays. Spin quantization frames orientation needs care. It must be the same for production and decay. We use KKMC routines to transfer $h^{i}_{\pm}$ to lab frame and another routines to transfer back to $\tau^{\pm}$ rest frame but oriented as in New Physics calculation. In this way reference frames are easier to control and impact of photons on phase space parametrisations is not of great complication. Note that the solution, if weights are calculated during KKMC run, is valid for all $\tau$ decays, while in case of TauSpinner, only for main $\tau$ decay channels spin effects are included.
@@
Figure 3: Ratio of number of events with and without weak dipole moments, in function of acoplanarity $\varphi$ at $\sqrt{s}=M_{Z}$. The selected events of scattering angles $\cos(\theta)<0$ are taken. The top left plot for ${\rm Re}(X)=0.0004$, the top right plot for ${\rm Re}(Y)=0.0004$, the bottom left for ${\rm Im}(X)=0.0004$, and the bottom right for ${\rm Im}(Y)=0.0004$ are taken . For the imaginary form-factors, additional constraint $E_{\pi^{+}}>E_{\bar{\nu}_{\tau}}$ is taken the $\tau^{+}$ side. The form-factors $A(M_{Z}^{2})=B(M_{Z}^{2})=0$ are set. The decays $\tau^{-}\to\pi^{-}\pi^{0}\nu$ and $\tau^{+}\to\pi^{+}\nu$ are taken. Plots taken from Ref. [9].
@@ 5 Conclusions and outlook
I have presented some new developments and and new applications for Monte Carlo programs family arranged around KKMC. Let me finally point to some possible, near future, activities. The big issue important for systematic errors are ambiguities of $\alpha_{QED}(s)$, the function of electromagnetic coupling. For the Monte Carlos, its values for $s$ starting from 00 up to beyond $m_{Z}^{2}$ are needed. Also the g-2 measurements revived interest in possible signatures with $\tau$ leptons: real and virtual. I have presented some work relevant for that. But what about related systematic ambiguities?
Can modified KKMC be helpful in such studies? For $\alpha_{QED}(s)$ dispersion relations and measurements of $e^{+}e^{-}\to\pi^{+}\pi^{-}\pi^{0}(\pi^{+}\pi^{-})$ at $\sqrt{s}\sim 1GeV^{2}$ is necessary. This energy range can be achieved with Belle 2 data and events with radiative return. The invariant mass of the hadronic final state can be small if it is accompanied with hard initial state bremsstrahlung photon. The question arise if KKMC can be adopted and become helpful.
@@
I have not presented further details on arrangements for $\tau$ decays and for software, except what was explained in introduction.

### v4

@@ Abstract
One of the purposes of High Energy accelerator experiments is confrontation of theory and measurements in ever new realms. Any new agreement extends theory applicability domain, any discrepancy hints to unexplained. That calls for better calculations or new, deeper theory. Often one has to search for small contributions over large Standard Model background. Multidimensional signatures and complex background subtraction cuts imply that Monte Carlo techniques are indispensable. My talk included description of: KKMC Monte Carlo for $e^{+}e^{-}\to\tau^{+}\tau^{-}(n\gamma)$ (with $\tau$ decays), generation of additional lepton pairs of SM and New Physics, and contributions from anomalous magnetic and electric dipole moments.
@@ 1 Introduction
My talk is devoted to KKMC Precision Monte Carlo for $e^{+}e^{-}\to\tau^{+}\tau^{-}(n\gamma)$ , with $\tau$ decays included, and its associated projects. The main design purposes, as in the past, include: (i) precision simulation of lepton pair production, (ii) flexible set up for initialization of user defined hadronic currents used in $\tau$ decays. There are new aspects which gain importance. The C++ version of KKMC is available, which has better than the FORTRAN predecessor arrangements for beam-spread for example. But the most important aspect was to preserve continuity of numerical tests, an essential ingredient of work see Ref. [1]. The long road toward enhanced precision of QED part started already  [2], but this road toward third order matrix elements is long and perilous. There is a need for new contributors to enter, to preserve skills and continuity of the projects, and at the same time to address new challenges.
Let us now turn attention to specific projects, to their limitations too. The C++ version of KKMC is public, but is not yet adopted to lower energies (the function for photon vacuum polarization $\Pi_{\gamma\gamma}(s)$ to be obtained from dispersion relations is clearly defined, but needs to be then adjusted). There are clear manpower issues: help in installations, set-ups and of some phenomenology applications is not as robust as it was in the past. For the F77 version as installed in Belle2 I take care and help, but main burden is now on somebody’s else shoulders.
Let us turn now attention to generation of additional lepton pairs originating from SM or New Physics interactions. Pair emission is a necessary step toward third order matrix element installation. From the algorithmic point of view, treated as radiative corrections it is incompatible with exponentiation, because the corresponding crude distribution does not feature conformal symmetry. This symmetry is essential in building a relation between crude level Monte Carlo with matrix elements of eikonal form. It is also important for matching line-shape distribution (intermediate resonances) and initial state bremsstrahlung. Alternative solutions are necessary. Real pair emission can be achieved by running of KKMC with appropriate flags for pair emission part of loop vertex correction switched on. Then a simultaneous run, of four fermion final state Monte Carlo program, like KORALW [3], is necessary. Alternatively, for final state pair emissions, Monte Carlo of the after-burn type, PHOTOS is available [4]. The program enables full phase-space coverage. Although tests with Matrix element simulations indicate that precision is sufficient for today applications, some studies of related systematic ambiguities are included in Ref. [5]. These publications do not match the FCC precision, and further work is needed.
@@
Let us now turn attention to more technical aspects. The process of migration from FORTRAN to C++ is advancing well. The C++ Versions of PHOTOS Monte Carlo for radiative corrections in decays [4] and KKMC [1] are already published. The TAUOLA Monte Carlo for $\tau$ decays is prepared for such transformation too, but actual implementation is not completed. Its internal structure was already prepared some time ago [10]. Further improvements are now in Belle 2 hands. Fits and evaluation of ambiguities of explored data are necessary. The program was prepared to run with user provided hadronic currents, which could be coded in C++ and activated with redefinition of pointers. That solution was temporarily but partly abandoned. Step back was due to requirements of Belle software organization.
My experience with PHOTOS where migration to C++ has been completed is that the best is to translate program with young researchers’ participation. They can observe that migration does not go too fast, because existing arrangements for future and for tests could be then easily lost. Participation of researchers with distinct profile is helpful, and prepares us for future takeover of the projects, independently whoever it may be.
The software for Belle 2 is working well, and I am involved personally. Important contributor is now Swagato Banerjee. Installation and initialization for other platforms pose now-days problems especially for C++ version of KKMC. For example proper loading of root library was bringing difficulties, and it was reported to us, but our response was slow. That is one of the most visible aspect of the man power issue for our community, but passing expertise to the next generation is potentially larger issue, even if related challenges are not imminent to see.
@@ 2 Basics of precision Monte Carlo
Let me stress at first KKMC is based on “matrix element $\times$ full and exact phase space”. Neither KKMC nor PHOTOS employ a shower-like algorithm. They are both non-Markovian. Generation starts with Poisson distribution in number of photon candidates. In case of KKMC tower of simplified/improvable structures starts from one-dimensional spectrum, then an internal generator exploiting eikonal matrix elements valid all over full multi-photon phase space is devised. It was a highly non-trivial step to achieve. Next step is exploiting bremsstrahlung part of QED matrix elements and Yennie–Frautschi–Suura exponentiation [11]. That is not all what is needed: the electroweak and strong interaction parts need to be taken in as well. Finally, light lepton pair emissions must be taken into account. I have omitted the complexity of presamplers necessary due to collinear configurations, which are naturally regulated by electron (or other lepton) masses. Finally, the library of Matrix Elements devised as input for event weight, sometimes called “model weight”, represents an independent module see Fig. 1. The term “model weight” may be misunderstood. There is no room for arbitrary modeling. Amplitudes in use are the result of formal fixed order matrix elements. For details see [12].
@@
Note that for phase space organization, in KKMC design, conformal symmetry of multi-photon phase space and eikonal parts of QED amplitudes is fundamental. That is why, algorithm is suitable for many applications, path for implementation of fixed order amplitudes is defined. At present, second order matrix elements are installed. In case of PHOTOS different solution is used for exact multi-photon phase space generation. It is restricted to final state radiations only, and solution is not explored beyond use of the first order matrix elements. On the other hand, emission of massive particles was easier to implement. For details seem [5, 4] and references therein.
@@ 3 Additional pairs
For pair emissions in final states modified algorithm of PHOTOS Monte Carlo can be used. In this case phase space is treated exactly, and no approximations are used, independently of whether the pair emission is based on QED calculated matrix element or of New Physics model. As effects of pair emissions are not large, matrix elements are simplified (in improvable way though)! Tests/comparisons with matrix element based calculations are essential part of the work. As an example Fig. 2 is provided.
@@ 4 Anomalous dipole moments
For the simulation of New Physics effects, simplified kinematic for implementation is usually sufficient. There can be two weights calculated, for the cross-section:
@@
The $R_{ij}$ depends on kinematic of $\tau$-pair production, $h^{i}_{\pm}$ on $\tau^{\pm}$ decays. Spin quantization frames orientation needs care. It must be the same for production and decay. We use KKMC routines to transfer $h^{i}_{\pm}$ to lab frame and another routines to transfer back to $\tau^{\pm}$ rest frame but oriented as in New Physics calculation. In this way reference frames are easier to control and impact of photons on phase space parametrization is not of great complication. Note that the solution, if weights are calculated during KKMC run, is valid for all $\tau$ decays, while in case of TauSpinner, only for main $\tau$ decay channels spin effects are included.
@@
Figure 3: Ratio of number of events with and without weak dipole moments, in function of acoplanarity $\varphi$ at $\sqrt{s}=M_{Z}$. The selected events of scattering angles $\cos(\theta)<0$ are taken. The top left plot for ${\rm Re}(X)=0.0004$, the top right plot for ${\rm Re}(Y)=0.0004$, the bottom left for ${\rm Im}(X)=0.0004$, and the bottom right for ${\rm Im}(Y)=0.0004$ are taken . For the imaginary form-factors, additional constraint $E_{\pi^{+}}>E_{\bar{\nu}_{\tau}}$ is taken in the $\tau^{+}$ side. The form-factors $A(M_{Z}^{2})=B(M_{Z}^{2})=0$ are set. The decays $\tau^{-}\to\pi^{-}\pi^{0}\nu$ and $\tau^{+}\to\pi^{+}\nu$ are taken. Plots taken from Ref. [9].
@@ 5 Conclusions and outlook
I have presented some new developments and and new applications for Monte Carlo programs family arranged around KKMC. Let me finally point to some possible, near future, activities. The big issue important for systematic errors are ambiguities of $\alpha_{QED}(s)$, the function of electromagnetic coupling. For the Monte Carlos, its values for $s$ starting from 00 up to beyond $m_{Z}^{2}$ are needed. Also, the g-2 measurements revived interest in possible signatures with $\tau$ leptons: for real productions or through loops effects. I have presented some work relevant for that. But what about related systematic ambiguities?
Can modified KKMC be helpful in such studies? For $\alpha_{QED}(s)$ dispersion relations and measurements of $e^{+}e^{-}\to\pi^{+}\pi^{-}\pi^{0}(\pi^{+}\pi^{-})$ at $\sqrt{s}\sim 1GeV^{2}$ is necessary. This energy range can be achieved with Belle 2 data and events with radiative return. The invariant mass of the hadronic final state can be small if it is accompanied by hard initial state bremsstrahlung photon. The question arise if KKMC can be adopted and become helpful.
@@
I have not presented further details on arrangements for $\tau$ decays and for software, except what was explained in the introduction.